Philip sold a certain number of athletic shoes for $2160. He sold 60% of them at $40 each and the rest at 25% discount. How many pairs of athletic shoes did sell at 25% discount?
Answer by Maths68(1474)

Let
Total pairs of shoes sold = x
60% of sold shoes = x*60%=x*60/100=x*3/5=3x/5
Cost of a pair of 60% sold shoes = $40
Total Cost of 60% of sold shoes = 40*3x/5=8*3x=24x
Cost of a pair sold shoes (Remaining) = 40-(40)*(25/100)= 40-(40)*(1/4)=40-10=$30
40% of sold shoes (Remaining) = x*40%=x*40/100=x*2/5=2x/5
Total Cost of 40% of sold shoes (Remaining) = 2x/5 * 30=2x*6=12x
Total cost of all shoes = Total Cost of 60% of sold shoes + Total Cost of 40% of sold shoes (Remaining)
2160=24x+12x
2160=36x
Divide by 36 both sides
2160/36=36x/36
60=x
x=60
60 pairs of shoes all together sold
40% of sold shoes (Remaining) = x*40%=x*40/100=x*2/5=2x/5=2(60)/5=2*12=24
60% of sold shoes = x*60%=x*60/100=x*3/5=3x/5=3(60)/5=36
24 pairs of shoes sold at 25% discount on per pair
36 pairs of shoes sold at $40 per pair
Check
Total cost = Total Cost of 60% of shoes sold at $40/pair + Total Cost of 40%
2160=36*40+24*30
2160=1440+720
2160=2160